OpenStudy (howard-wolowitz):

A local video store owner is examining his two stores, the one on 29th Avenue and the other on Division. The 29th store does 70% of the dealers business and the Division store does the rest. In the Division store, 40% of all rentals are DVDs. In the 29th store, 30% of all rentals are DVDs. If a customer is selected at random, what is the probability that she will rent a DVD?

OpenStudy (knov):

costumer ?

OpenStudy (howard-wolowitz):

I fixed it

OpenStudy (knov):

can you ss a question written in your course book ?

OpenStudy (howard-wolowitz):

are you saying this is a course book question?

OpenStudy (knov):

what i mean is the question is confusing. Can you send me the link of the source or ss.

OpenStudy (knov):

can you do that ?

OpenStudy (howard-wolowitz):

no, because this is a homework question so there is only my sheet of paper

OpenStudy (knov):

i see. Sorry can't help you with it.

OpenStudy (howard-wolowitz):

no problem

hartnn (hartnn):

have you tried? know how to start?

OpenStudy (howard-wolowitz):

work is this: 70% in roomA, and 30% in roomB ------------------------------- we ask the 700 people in roomA if they got a DVD, and 30% of them say yes we ask the 300 people in roomB if they got a DVD, and 40% of them say yes -------------------------------- so, out of 1000 people, how many are there altogether that have a DVD? 7*30, and 3*40 .. 210+120 = 330 --------------------------------- 330 people out of 1000 have a DVD

OpenStudy (howard-wolowitz):

0.33

hartnn (hartnn):

lets verify it using my way. she will rent the DVD when she chooses the 29th store(70%) AND chooses a DVD (30%) OR she chooses Division store(30%) AND she chooses a DVD (40%) = 0.7*0.3 + 0.3*0.4 = 0.3*1.1 = 0.33 good!

hartnn (hartnn):

AND -- multiplication, OR -- addition

OpenStudy (howard-wolowitz):

ok thank you

hartnn (hartnn):

welcome ^_^